Apply the method of variations of parameters to solve the following differential equations:
a) x^2y double prime+xyprime-y=x^2e^x
1
Expert's answer
2020-03-18T18:22:35-0400

"x^2y''+xy'-y=x^2e^x"

We will replace

"x=e^t,\\
t=ln|x|,\\
y'=\frac{dy}{dx}=\frac{y'_t}{x'_t}=\frac{y'_t}{e^t}=e^{-t}y'_t,\\
y''=\frac{dy'}{dx}=\frac{\frac{dy'}{dt}}{\frac{dx}{dt}}=\frac{y''_te^{-t}-e^{-t}y'_t}{e^t}=e^{-2t}(y''_t-y'_t)\\"

Input in equation

"e^{2t}e^{-2t}(y''_t-y'_t)+e^te^{-t}y'_t-y=e^{2t}e^{e^t}"

get the equation

"y''-y=e^{2t}e^{e^t}"

Consider homogenous equation

"y''-y=0\\
\lambda^2-1=0\implies\lambda_1=1, \lambda_2=-1\\
y_1=e^t, y_2=e^{-t}\\
y_0=c_1e^t+c_2e^{-t}"

Solution of equation find in form

"y=c_1(t)e^t+c_2(t)e^{-t}"

"c_1, c_2" find from the system

"\left\{
\begin{matrix}
 c_1'(t)e^t+c_2'(t)e^{-t}=0 \\
 c_1'(t)e^t-c_2'(t)e^{-t}=e^{2t}e^{e^t}
\end{matrix}
\right.\\
\left\{
\begin{matrix}
 c_1'(t)=\frac{1}{2}e^{t}e^{e^t}\\
 c_2'(t)=-\frac{1}{2}e^{3t}e^{e^t}
\end{matrix}
\right.\\
c_1(t)=\int \frac{1}{2}e^{t}e^{e^t}dt=\frac{1}{2}e^{e^t}+k_1\\
c_2(t)=-\int \frac{1}{2}e^{3t}e^{e^t}dt=\\
=-\frac{1}{2}(e^{2t}e^{e^t}-2e^{t}e^{e^t}+2e^{e^t})+k_2\\
y=(\frac{1}{2}e^{e^t}+k_1)e^t+\\
+(-\frac{1}{2}(e^{2t}e^{e^t}-2e^{t}e^{e^t}+2e^{e^t})+k_2)e^{-t}"

Than

"y=(\frac{1}{2}e^x+k_1)x+(-\frac{1}{2}(x^2e^x-2xe^x+2e^x)+k_2)x^{-1}"